Our teacher didn't explain much about using Cramer rule with fractions so im at a lost at this equation.


{(2/3)x + (1/5)y = 3
{(3/2)x + (5/3)y = (2/3)


If it makes it any easier, our teacher said we can also use subtitution and elimination method.

Answer by stanbon(57246)

Our teacher didn't explain much about using Cramer rule with fractions so im at a lost at this equation.
{(2/3)x + (1/5)y = 3
{(3/2)x + (5/3)y = (2/3)
If it makes it any easier, our teacher said we can also use subtitution and elimination method.
--------------------
The fact that you have fractions does not make any difference in
using Cramer's Method.
But you can get rid of the fractions:
Multiply thru the 1st equation by 15 to get:
10x + 3y = 45
Multiply thru the 2nd equation by 6 to get:
9x + 10y = 4
------------------
Now, use Cramer or elimination to get:
x = 6
y = -5
==========
Cheers,
Stan H.